Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below Indicate the missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct labor cost and the rate is the same. (Round overhead rate to 2 decimal places, eg. 15.25 and final answers to O decimal places, e.g. 5,275.) Case A Case B Case C Direct materials used Direct labor Manufacturing overhead 55450 (a) $ 91,800 $73,650 60.000 144,800 86880 (d) 323480 (e) 22,100 345580 (f) 14,100 331480(g) 36,000 applied Total manufacturing costs Work in process 1/1/17 Total cost of work in process Work in process 12/31/17 Cost of goods manufactured 151,450 56450 (b) 207,900 212,850 18,400 0) 9300 (c) 198,600 227,000
Answer #1

Calculation of Direct labor(h) for case C:

Direct material used + Direct labor + Manufacturing overhead(i.e to be applied on the basis of direct labor cost) = Total manufacturing costs

$73650 + h + (h*60%) = 212850

Therefore h = $87,000

Calculation of manufacturing overhead(i) for case C:

Direct material used + Direct labor + Manufacturing overhead = Total manufacturing costs

$73650+h+i = 212850

$73650+$87000+i = 212850

Therefore i = $52,200

Calculation of total cost of work in progress(j) for case C:

Total manufacturing costs + work in progress 1/1/17 = Total cost of work in progress

$212850 + $18400 = j

Therefore j = $2,31,250

Calculation of work in progress 12/31/17 for case C:

Total cost of work in progress - work in progress 12/31/17 = Cost of goods manufactured

$2,31,250 - k = $2,27,000

k = $4,250

CaseA Case B Case C $91,800 144,800 86,880(d) 323,480(e) 22,100 345,580(f) 14100 $73,650 87,000(h) 52,200(i) Direct materials used 55,450(a) Direct labor 60,000 Manufacturing overhead applied Total manufacturing costs Work in progress 1/1/17 Total cost of work in progress work in progress 12/31/17 Cost of goods manufactured 36,000 151,450 212,850 56450(b) 18,400 231,250(j) 4250(k) 207,900 9300(c) 198,600 331,480(g) 227,000